OpenStudy (anonymous):

Which of the following is the formula for this geometric sequence? 1, -3, 9, -27,... Answer an = -1(3)n-1 an = 1(-3)n-1 an = 3(-1)n-1

OpenStudy (anonymous):

\[ a_n=a_1\times \left(\frac{a_2}{a_1}\right)^{n-1} \]

OpenStudy (anonymous):

\(a_1=1,a_2=-3\)